1. The average accounting rate of return (AAR):

A) considers the time value of money.

B) measures net income as a percentage of the sales generated by a project.

C) is the best method of financially analyzing mutually exclusive projects.

D) is the primary methodology used in analyzing independent projects.

E) is similar to the return on assets ratio.

2. Consider the following information:

State of

Economy

Probability of

State of Economy

Rate of Return

if State Occurs

Stock A

Stock B

Recession

.04

.097

.102

Normal

.72

.114

.133

Boom

.24

.156

.148

The market risk premium is 7.4 percent, and the risk-free rate is 3.1 percent. The beta of Stock A is ________ and the beta of Stock B is ________.

A) 1.25; 1.89

B) 1.47; 1.76

C) 1.21; 1.76

D) 1.47; 1.41

E) 1.25; 1.41

3. The common stock of Manchester & Moore is expected to earn 14 percent in a recession, 7 percent in a normal economy, and lose 4 percent in a booming economy. The probability of a boom is 15 percent while the probability of a recession is 5 percent. What is the expected rate of return on this stock?

A) 8.5 percent

B) 8.7 percent

C) 5.7 percent

D) 7.5 percent

E) 6.2 percent

4. Assume the spot rate on the pound is £.5920 and the 6-month forward rate is £.5929. Also assume interest rate parity holds and the current six-month risk-free rate in the United States is 3.1 percent. What must the six-month risk-free rate be in Great Britain?

A) 3.25 percent

B) 2.87 percent

C) 2.94 percent

D) 3.10 percent

E) 3.52 percent
Answer #1

1. Average accounting rate of return = Average net annual profit/ initial investment

It does not include time value of money It measures net income as a percentage of investment. It is not the bes method to analyse mutually exclusive projects and not the primary methodology to analyse independent project. However since investment is used to create assets , this formula is similar to return on assets ratio. (option E)

2.  The formula for Expected return is

  Expected Return = { p; * R

RA = 0.04*0.097 +0.72*.114+0.24*0.156

=.1234 = 12.34%

RB = 0.04*0.102 +0.72*.133+0.24*0.148

=.1354 = 13.54%

by CAPM,

Return of a stock = Risk free rate + beta of stock * market risk premium

So, 12.34% = 3.1%+ Beta (A) * 7.4%

=> Beta (A) = 9.24%/7.4% = 1.25

Similarly, for stock B

13.54% = 3.1% +Beta (B) * 7.4%

=> Beta (B) = 10.44%7.4% = 1.41

So option E is correct

3. The probability of normal economy = 1- probability of boom - probabiity of recession = 1-0.05-0.15 =0.8

The formula for Expected return is

  Expected Return = { p; * R

So Expected return of Manchester & Moore is

R = 0.05*7%+0.8*7% +0.15*(-4%)

=5.7% (option C)

4. From interest rate parity

Forward rate after 6 months/Spot rate  = (1+6 month risk free rate of Great Britain)/(1+ 6 month risk free rate of USA)

=> 0.5929/0.5920 = (1+6 month risk free rate of Great Britain) / (1+0.031)

=> (1+6 month risk free rate of Great Britain) = 1.00152 * 1.031 = 1.0325

=> 6 month risk free rate of Great Britain = 0.0325 = 3.25% (option A)